OpenStudy (anonymous):

complex fractions- 17and 3/5 over 11 how would i solve

OpenStudy (anonymous):

17 and 3/5 = (17*5+3)/5=88/5

OpenStudy (anonymous):

\[\frac{\frac{88}{5}}{\frac{11}{1}}=\frac{88}{5}*\frac{1}{11}=\frac{8*11}{5*11}=\frac{8}{5}\]

OpenStudy (anonymous):

8/5 = 1 and 3/5